What is the probability of the existence of aliens?
Well the universe in infinitely large with plenty of other solar systems, so to believe that life started on some other planet than ours somewhere out there isn't wrong. Whether they are more advanced then us is another thought.
Actually, although it may seem like it, it is not 50-50. The answer's more like 55-45 This is because the 'head' side is heavier than the 'tails' side. So it is more likely to be tails 3 heads 1
There are 12 face cards in a standard deck of 52 cards. The odds of the first card being a face card is 12/52. If the first card drawn is a face card then there are 11 face cards remaining in the deck of 51 cards. The odds of a second draw of a face card is then 11/51. If both the first two cards drawn were face cards then the deck has 10 face cards in 50 total card. The odds of the third card also being a face card is 10/50. The total probability is (12/52)*(11/51)*(10/50) = 0.009954751 or just under one percent of the time.
If something is certain the probability is?
The probability of something that is certain is 100% The probability of something that is completely impossible is 0%
If a coin is tossed two times what is the probability of getting tails two times in a row?
25% or 0.25 Probability of one tail is 0.50. Since two tails are independent events, the probability is 0.5 x 0.5 = 0.25
What is addition and multiplication theorems on probability with examples?
When considering the probability of two different events or outcomes, it is essential to clarify whether they
are mutually exclusive or independent. If the events are mutually exclusive, then the probability that either
one or the other will occur equals the sum of their individual probabilities. This is known as the law of
addition.
If, however, two or more events or outcomes are independent, then the probability that both the first and
the second will occur equals the product of their individual probabilities. This is known as the law of
multiplication.
The probability of heads on each of the coins is 1/2 or 0.50, and the die (we will assume a fair die, numbered 1 to 6) coming up 4 is 1/6. So we have: 1/2 x 1/2 x 1/6 = 1/24 = 0.04167 or approximately 4.2% chance of these three events occurring.
How is the accuracy of a poll expressed?
sampling error
Another VIew: PROBABILITY - as expressed in plus or minus a certain number of percentage points.
True
This is correct. For example the probability of tossing a coin so that it comes up heads is 1/2 and the probability that it comes up tails is also 1/2. The probability that it will come up either heads or tails is 1.
What is the probability of rolling a number less than 5 or a odd number?
You didn't mention whether this was one die or two dice. If one die, then the possibilites are 1 and 3, so this is 2 outcomes out of 6 or 1/3. Probability can stated as 0.33 or 33%. If two die, there is 2 possibilities: (1,2) and (2,1) out of 36 outcomes which is 1/18 = 0.055. Probability can be stated as 0.055 or 5.5%.
What is the probability density of the wave function?
See related link, In quantum mechanics, a probability amplitude is a complex number whose modulus squared represents a probability or probability density. For example, the values taken by a normalised wave function ψ are amplitudes, since |ψ(x)|2 gives the probability density at position x. Probability amplitudes may also correspond to probabilities of discrete outcomes.
What is probability of obtaining a number different from 11 when a single fair die is tossed?
A single fair die has the numbers 1 to 6, so when a single fair die is tossed the
probability of obtaining a number different than 11 is: P(x diff than11) = 1.
How do you find the experimental probability that a particular result will occur?
You carry out the experiment a large number of times. Count the number of times it was carried out (n). Count the number of times in which the particular outcome occurred (x). Then, the experimental probability for that even is x/n.
There are 12 months to choose from There are 7 months with 31 days in them.
The probability of choosing a 31-day month is 7/12.
There are 30 students total. 10 of them have brown hair. 10/30 equals 1/3. Answer: 1/3
What is the third question in the four-way test?
The questions on the four-way test are:
What are three events in which the probability is 1?
If an event has a probability of 1, it will happen no matter what.
The probability of rolling a number x, such that 1 ≤ x ≤ 6, on a standard 6 sided die, is 1.
The probability of the temperature being > absolute 0, is 1.
With a standard 52 card deck of card, P(drawing a spade or drawing a club or drawing a heard or drawing a diamond) is 1.
In these situations, there is no variable that can affect the event. It will happen no matter what.
24 out of 6497400 = 1 out of 270725.
What is the probability of choosing a picture card from a standard pack of playing cards?
There are fifty-two cards in a deck. Twelve are picture cards (king, queen, jack for each of four suits.) So your chances of getting a picture card are twelve out of fiftytwo, which you can reduce down to three out of thirteen.
What is the probability that a family of 4 children has exactly 3 boys?
0.25 binomial distribution, where p=0.5, q=0.5, x=3, n=4 4!/(3!*1!)*0.530.51 = 0.25 Also can be solved by identifying each event possible and related probability. There are 4 ways this can occur (first child is a girl, second child is a girl, third child is a girl and fourth child is a girl) and there is a 0.54 chance of each of these events occurring. Prob= 4 *0.0625 = 0.25
The chance of picking a heart is 1/4 as there are four suits. The chance of picking an ace is 1/13, as there are 4 aces in a deck of 52. If I want the probability of drawing a heart of an ace, I can add these probabilities together, but I have double counted on possibility, the ace of hearts. So: Probability = 4/52 + 13/52 - 1/52 = 16/52 or 4/13 Probability = 0.308 expressed as a fraction.